# Sensory Data Sharing in ISAC System: Fundamental Limits and Waveform Design

### v2

@@ II-B The Estimate and Compress Strategy
In the EC strategy, the encoder transmits the estimate $\mathsf{\tilde{S}}$ extracted from the observation $\mathsf{Z}_{i}$, instead of transmitting $\mathsf{Z}$ itself. Therefore, the coding scheme may be recast by
@@
Although the estimator and encoder are defined in a block form for generality, the following Lemma 1 [14] show that the optimal estimator $h$ can be achieved by a symbol-by-symbol estimation.
@@
which is independent to the choice of encoder and decoder.
The detailed proof can be found in [14]. By applying the optimal estimator (4), the estimate $\mathsf{\tilde{S}}_{i}$ only depends on $\mathsf{X}_{i}$ and $\mathsf{Z}_{i}$ but independent of past inputs and outputs. Thus, the encoder of this scenario is reduced to a symbolwise encoding process. Moreover, by introducing the estimate $\mathsf{\tilde{S}}_{i}$, the CAS distortion $D$ can be divided into two parts, i.e., sensing distortion $D_{s}$ between $\mathsf{S}$ and $\mathsf{\tilde{S}}$ and communication distortion $D_{c}$ between $\mathsf{\tilde{S}}$ and $\mathsf{\hat{S}}$. Before presenting the MSST, we will give the definitions of $D_{s}$ and $D_{c}$.
In the sensing process, the estimation-cost function with respect to (w.r.t.) the ISAC channel input realization $\mathsf{X}=x$ can be defined by
@@
Thus, a given sensing distortion $D_{s}$ satisfies
$\mathbb{E}\left[d(\mathsf{S},\mathsf{\tilde{S}})\right]=\mathbb{E}\left[e(\mathsf{X})\right]\leq D_{s}.$ (6)
In the communication process, the channel capacity depends on the ISAC channel input $\mathsf{X}$ which is constrained by a desired sensing distortion $D_{s}$ in the sensing process. This strong coupling property leads to a performance tradeoff between the sensing distortion and the achievable communication channel capacity. To proceed, we give the following definition.
Definition 1: The constrained channel capacity with sensing distortion $D_{s}$ and resource cost $B$ can be defined by
@@
The sequence $\{\mathsf{X}_{i}\}_{i\geq 1}$ is i.i.d. with distribution $P_{\mathsf{X}}(x)$ for achieving the capacity-distortion-cost region. By recalling the i.i.d. state sequence $\{\mathsf{S}_{i}\}_{i\geq 1}$, the observation $\{\mathsf{Z}_{i}\}_{i\geq 1}$ is also i.i.d. subject to the distribution
$P_{\mathsf{Z}}(z)=\sum_{s\in\mathcal{S}}\sum_{x\in\mathcal{X}}Q_{\mathsf{Z|XS}}(z|x,s)P_{\mathsf{S}}(s)P_{\mathsf{X}}(x).$ (10)
Furthermore, by adopting the symbol-by-symbol optimal estimator (4), the estimate $\{\tilde{\mathsf{S}}_{i}\}_{i\geq 1}$ is i.i.d. with the following distribution
@@
with $\mathbb{I}(\cdot)$ being the indicator function. Therefore, for the i.i.d. source $\mathsf{\tilde{S}}$, the relationship between $D_{c}$ and bit rate $R$ can be characterized by the rate-distortion theory.
@@
$R_{\text{IT}}(D_{c})=\mathop{\min}\limits_{P_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s}):\mathbb{E}\left[d(\mathsf{\tilde{S}},\mathsf{\hat{S}})\right]\leq D_{c}}I(\mathsf{\tilde{S}};\mathsf{\hat{S}}),$ (12)
where the minimization is over all conditional distributions $P_{\mathsf{\mathsf{\hat{S}}|\mathsf{\tilde{S}}}}(\hat{s}|\tilde{s})$ for which the joint distribution $P_{\mathsf{\tilde{S}\hat{S}}}(\tilde{s},\hat{s})=P_{\mathsf{\tilde{\mathsf{S}}}}(\tilde{\mathsf{s}})P_{\mathsf{\mathsf{\hat{S}}|\mathsf{\tilde{S}}}}(\hat{s}|\tilde{s})$ satisfies the expected distortion constraint. The rate-distortion function $R_{\text{IT}}(D_{c})$ demonstrates the minimal rates required to achieve a given communication distortion $D_{c}$.
@@ III Modified Source-Channel Separation Theorem
In this section, we develop the modified source-channel separation theorem tailored for the dual-functional signaling CAS systems, aiming to elucidate the relationships between the CAS distortion, coding rate and communication channel capacity. The MSST retains a form analogous to the conventional source-channel separation theorem with distortion in lossy data transmission [28, Theorem 10.4.1].
@@
$\mathbb{E}\left[d(\mathsf{S},\mathsf{\hat{S}})\right]=\mathbb{E}\left[d(\mathsf{S},\mathsf{\tilde{S}})\right]+\mathbb{E}\left[d(\tilde{\mathsf{S}},\mathsf{\hat{S}})\right],$ (13)
implying that the CAS distortion $D$ is equal to $D_{s}+D_{c}$.
Note that the separability condition in Definition 3 appears to be strong. Fortunately, we demonstrate that the mean squared error (MSE) or quadratic distortion, a widely used distortion metric in parameter estimation, satisfies the separability condition. Specifically, we have
$D$ $=\mathbb{E}\left[\left\lVert\mathsf{S}-\hat{\mathsf{S}}\right\rVert_{2}^{2}\right]$ (14)
$\mathop{=}\limits^{(a)}\mathbb{E}\left[\left\|\mathsf{S}-\tilde{\mathsf{S}}\right\|_{2}^{2}\right]+\mathbb{E}\left[\left\|\tilde{\mathsf{S}}-\hat{\mathsf{S}}\right\|_{2}^{2}\right]\mathop{=}\limits^{\Delta}D_{s}+D_{c},$
where $(a)$ holds from the properties of the conditional expectation [20, Appendix A] with $\mathsf{\tilde{S}}=\mathbb{E}\left[\mathsf{S}|\mathsf{Z},\mathsf{X}\right]$ being the optimal estimator  [^1] We highlight that the (4) is exactly the minimum MSE (MMSE) estimator. An arbitrary non-optimal estimator except MMSE, which cannot apply the conditional expectation properties, may not satisfy the separability condition. . Now, we are ready to propose the MSST.
Theorem 1: The CAS distortion $D$ is achieved within a separable distortion metric, if and only if
@@
where $C_{\text{IT}}(D_{s},B)$ and $R_{\text{IT}}(D_{c})$ are the constrained channel capacity (7) and the rate-distortion function (12), respectively.
@@ III-A Converse
We start with a converse to show that any achievable coding scheme must satisfy (15). Consider a $(2^{nR},n)$ coding scheme defined by the encoding and decoding functions $\phi$ and $\psi$. Let $\mathsf{\tilde{S}}^{n}=h^{\star}(\mathsf{X}^{n},\mathsf{Z}^{n})$ be the estimate sequence as given in (3) and $\mathsf{\hat{S}}^{n}=\psi(\phi(\mathsf{\tilde{S}}^{n}))$ be the reconstruction sequence corresponding to $\mathsf{\tilde{S}}^{n}$.
@@ III-B Achievability
1) Codebook Generation: In source coding with rate distortion code, randomly generate a codebook $\mathcal{C}_{s}$ consisting of $2^{nR}$ sequences $\mathsf{\hat{S}}^{n}$ which is drawn i.i.d. $\sim P_{\mathsf{\hat{S}}}(\hat{s})$. The probability distribution is calculated by $P_{\mathsf{\hat{S}}}(\hat{s})=\sum_{\tilde{s}}P_{\mathsf{\tilde{S}}}(\tilde{s})Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$, where $P_{\mathsf{\tilde{S}}}(\tilde{s})$ is defined in (11) and $Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$ achieves the equality in (12). In channel coding, randomly generate a codebook $\mathcal{C}_{c}$ consisting of $2^{nR}$ sequences $\mathsf{X}^{n}$ which is drawn i.i.d. $\sim P_{\mathsf{X}}(x)$. The $P_{\mathsf{X}}(x)$ is chosen by satisfying the constrained capacity with estimation- and resource-cost in (7). Index the codeword $\mathsf{\tilde{S}}^{n}$ and $\mathsf{X}^{n}$ by $w\in\{1,2,\cdots,2^{nR}\}$.
@@
$(\mathsf{\tilde{S}}^{n},\mathsf{\hat{S}}^{n}(w))\in\mathcal{T}^{(n)}_{d,\epsilon_{s}}(P_{\mathsf{\tilde{S}\hat{S}}}(\tilde{s},\hat{s})),$ (18)
where $\mathcal{T}^{(n)}_{d,\epsilon_{s}}(P_{\mathsf{\tilde{S}\hat{S}}}(\tilde{s},\hat{s}))$ represents the distortion typical set [28] with joint probability distribution $P_{\mathsf{\tilde{S}\hat{S}}}(\tilde{s},\hat{s})=P_{\mathsf{\tilde{S}}}(\tilde{s})Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$. To send the message $w$, the encoder transmits $x^{n}(w)$.
@@
where $\mathcal{T}^{(n)}_{\epsilon_{c}}(P_{\mathsf{XY}}(x,y))$ represents the typical set with joint probability distribution $P_{\mathsf{XY}}(x,y)=P_{\mathsf{X}}(x)Q_{\mathsf{Y|X}}(y|x)$. If there exists such $\hat{w}$, it declares $\mathsf{\hat{S}}^{n}=\hat{s}^{n}(\hat{w})$. Otherwise, it declares an error.
@@
According to the achievability of lossy source coding [28, Theorem 10.2.1], we have $P_{e_{s}}$ tends to zero for sufficiently large $n$ whenever $R\geq R_{\text{IT}}(D_{c})$.
@@
Note that $(s^{n},x^{n}(w),\tilde{s}^{n})\in\mathcal{T}^{(n)}_{\epsilon_{e}}(P_{\mathsf{SX\tilde{S}}}(s,x,\tilde{s}))$ where $P_{\mathsf{SX\tilde{S}}}(s,x,\tilde{s})$ denotes the joint marginal distribution of $P_{\mathsf{SXZ\tilde{S}}}(s,x,z,\tilde{s})=P_{\mathsf{S}}(s)P_{\mathsf{X}}(x)Q_{\mathsf{Z|SX}}(z|s,x)\mathbb{I}\{\tilde{s}=\tilde{s}^{\star}(x,z)\}$, we have
@@
according to the typical average lemma [14]. In summary, the CAS distortion can be attained by
@@
where $(a)$ follows the condition of the separable distortion metric in MSST and we omit the terms containing the product of $P_{e_{s}}$ and $P_{e_{c}}$ in step $(b)$. Consequently, taking $n\to\infty$ and $P_{e_{s}},P_{e_{c}},\epsilon_{c},\epsilon_{e},\epsilon_{s}\to 0$, we can conclude that the expected CAS distortion (averaged over the random codebooks, state and channel noise) tends to be $D\to D_{c}+D_{s}$ whenever $R_{\text{IT}}(D_{c})\leq R\leq C_{\text{IT}}(D_{s},B)$.
@@
Remark: The sensing process determines the accuracy of the state information acquisition at the encoder, whereas the communication process governs the quantity of information transmitted to the decoder. Overemphasis on either process can result in substantial performance degradation. For instance, achieving optimal sensing performance alone may not ensure the accurate reconstruction at the decoder due to limitations in the communication channel capacity. Conversely, excessive communication capacity may be wasted if the sensory data lacks sufficient accuracy. This underscores the need to investigate optimal channel input as well as the waveform design that minimize the CAS distortion.
IV Channel Input Distribution Optimization for the CAS Systems
In this section, we explore the optimal input distribution for the CAS system under scalar channels, aiming to minimize the CAS distortion under the MMST constraint. Hereinafter, we adopt the MSE as the distortion metric to guarantee the MSST. Meanwhile, we drop the subscript “IT” in the MSST.
@@ IV-A Problem Formulation
$\mathcal{P}_{0}\left\{\kern 2.0pt\begin{aligned} \mathop{\text{min}}\limits_{P_{\mathsf{X}}(x)}\kern 5.0pt&D=D_{c}+D_{s}\\ \text{subject to}\kern 5.0pt&R(D_{c})\leq I(D_{s},B).\end{aligned}\right.$ (23)
By substituting the associated expressions into (23), problem $\mathcal{P}_{0}$ can be reformulated by
@@
$\mathbb{E}\left[d(\mathsf{S},\mathsf{\tilde{S}})\right]$ $=\mathbb{E}\left[\mathbb{E}\left[d(\mathsf{S},\tilde{s}(\mathsf{X},\mathsf{Z}))|\mathsf{X},\mathsf{Z}\right]\right]$ (25)
@@
where the sensing cost $e(x)$ defined in (5) can be expressed by
@@
$\mathbb{E}[d(\mathsf{\tilde{S},\hat{S}})]=\sum_{\tilde{s}}P_{\mathsf{\tilde{S}}}(\tilde{s})\sum_{\hat{s}}Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})d(\tilde{s},\hat{s}).$ (29)
@@
Here, we emphasize the unique challenge in solving problem $\mathcal{P}_{1}$. In the conventional BA algorithm, the primary approach involves constructing the Lagrangian function w.r.t. the variable and obtaining the optimal solution based on the first-order necessary condition. However, the probability distribution $P_{\mathsf{\tilde{S}}}(\tilde{s})$ in (29) and (30) is indeed a function of the variable $P_{\mathsf{X}}(x)$ as describe in (11). Consequently, deriving a closed-form expression of $P_{\mathsf{X}}(x)$ from first-order necessary condition is challenging, which imposes a significant obstacle to solving problem $\mathcal{P}_{1}$ by using the conventional BA algorithm. To address this issue, we propose a two-step BA-based optimal search method to seek for a sub-optimal solution in the following subsection.
@@ IV-B Two-step BA-based Optimal Search Algorithm
2) Sub-problem 2: We calculate the minimum communication distortion $D_{c}$ for the given channel input distribution $P_{\mathsf{X}}(x)$ and determine the communication distortion $D_{c}$, thereby obtaining the CAS distortion $D$.
By varying the preset sensing distortions, one may generate a set $\mathcal{D}$ collecting the CAS distortions calculated through the above two steps. Thus, the minimum CAS distortion and its corresponding distribution $P_{\mathsf{X}}(x)$ can be identified by the minimum value search. The detailed procedure is as follows.
$\bullet$ Constrained communication channel Capacity for a given sensing distortion.
@@
$\mathop{\text{max}}\limits_{P_{\mathsf{X}}(x)}\kern 10.0pt$ $I(\mathsf{X};\mathsf{Y})$ (31)
@@
$\mathop{\text{max}}\limits_{P_{\mathsf{X}}(x)}\kern 5.0pt$ $I(P_{\mathsf{X}}(x),Q_{\mathsf{Y|X}}(y|x))-\mu\sum_{x}P_{\mathsf{X}}(x)e(x)$ (32)
@@
In formula (34), the Lagrangian multiplier $\lambda$ and the posterior distribution $Q_{\mathsf{X|Y}}(x|y)$ are unknown. The former may be determined by a bisection search method such that $\sum_{x}P_{\mathsf{X}}(x)b(x)\to B$. For the latter, we have the optimal posterior distribution which maximizes the MI with given $P_{\mathsf{X}}(x)$ and $Q_{\mathsf{Y|X}}(y|x)$, expressed by [29, Lemma 9.1]

### v3

@@ II-B The Estimate and Compress Strategy
In estimate and compress strategy, the encoder transmits estimate $\mathsf{\tilde{S}}$ which is extracted from observation $\mathsf{Z}$, instead of transmitting $\mathsf{Z}$ itself. The coding scheme may be recast by
@@
Although the estimator and encoder are defined in a block form for generality, the following Lemma 1 [20] show that the optimal estimator $h$ can be achieved through a symbol-by-symbol estimation.
@@
which is independent to the choice of encoder and decoder. The detailed proof can be found in [20]. By applying the optimal estimator (4), the estimate $\mathsf{\tilde{S}}_{i}$ only depends on $\mathsf{X}_{i}$ and $\mathsf{Z}_{i}$ but independent of past inputs and outputs. Thus, the optimal encoder is reduced to a symbol-wise encoding process.
By applying the estimate and compress strategy, the information flow of the SDAS can be divided into two parts: 1) Obtain the estimate $\mathsf{\tilde{S}}_{i}$ from $\mathsf{Z}_{i}$ in sensing process; 2) Transmit the estimate $\mathsf{\tilde{S}}_{i}$ and recover $\mathsf{\hat{S}}_{i}$ from $\mathsf{Y}_{i}$ in communication process. Let us define the S&C distortions as the expectations of the estimate error and communication error, i.e.,
[…]
Next, we will show how S&C distortions can be achievable.
Let us define the estimate error function in sensing process with respect to (w.r.t.) the channel input realization $\mathsf{X}=x$ as
@@
Thus, the sensing distortion $D_{s}$ is achievable when $\mathbb{E}\left[e(\mathsf{X})\right]\leq D_{s}$ holds. Nevertheless, communication distortion $D_{c}$ strictly depends on channel capacity which in turn relies on the channel input $\mathsf{X}$. Thus, the strong coupling property, stemming from the shared channel input $\mathsf{X}$, leads to an inherent performance tradeoff between the S&C distortions. To quantify such relationship, we define the communication channel capacity for a given sensing distortion as follows.
Definition 1: The communication channel capacity constrained by sensing distortion $D_{s}$ and resource cost $B$ can be defined by
@@
The capacity restricts the maximum achievable rate $R$ of the communication channel. It is well-known that the relationship between rate $R$ and distortion $D_{c}$ can be characterized by the rate-distortion theory. Before presenting the rate-distortion function, we first show the communication source $\{\tilde{\mathsf{S}}_{i}\}_{i\geq 1}$ is i.i.d. with the distribution
@@
with $\mathbb{I}(\cdot)$ being the indicator function.
Note that the sequence $\{\mathsf{X}_{i}\}_{i\geq 1}$ is i.i.d. with distribution $P_{\mathsf{X}}(x)$ for achieving the capacity-distortion-cost region. Thus, the observation $\{\mathsf{Z}_{i}\}_{i\geq 1}$ is also i.i.d. due to the i.i.d. target’s state $\{\mathsf{S}_{i}\}_{i\geq 1}$. By adopting the symbol-by-symbol optimal estimator (4), we can immediately obtain (10). Therefore, for i.i.d. source $\mathsf{\tilde{S}}$, the rate-distortion function for the communication process can be defined as follows.
@@
[…]
According to the conventional source-channel separation theorem with distortion in lossy data transmission [32, Theorem 10.4.1], we have the conclusion that the communication distortion $D_{c}$ is achievable if and only if $R_{\text{IT}}(D_{c})\leq C_{\text{IT}}(D_{s},B)$. Building upon this, the remaining question is how to define the SDAS distortion in terms of the S&C distortions.
@@ III Modified Source-Channel Separation Theorem
In this section, we develop the modified source-channel separation theorem tailored for the SDAS systems, aiming to elucidate the relationships between the SDAS distortion, coding rate and channel capacity. In the MSST, we restrict the distortion metric satisfying the following separable condition.
@@
[…]
implying that the SDAS distortion $D$ is equal to $D_{s}+D_{c}$.
Intuitively, the separable condition does not necessarily hold in most scenarios. Fortunately, we demonstrate that the mean squared error (MSE) or quadratic distortion, a widely used distortion metric in parameter estimation, satisfying
$\mathbb{E}\left[\left\lVert\mathsf{S}-\hat{\mathsf{S}}\right\rVert_{2}^{2}\right]\mathop{=}\limits^{(a)}\mathbb{E}\left[\left\|\mathsf{S}-\tilde{\mathsf{S}}\right\|_{2}^{2}\right]+\mathbb{E}\left[\left\|\tilde{\mathsf{S}}-\hat{\mathsf{S}}\right\|_{2}^{2}\right],$ (13)
where $(a)$ holds from the properties of the conditional expectation [28, Appendix A] with $\mathsf{\tilde{S}}=\mathbb{E}\left[\mathsf{S}|\mathsf{Z},\mathsf{X}\right]$ being the optimal estimator in (4). Now, we are ready to propose MSST  [^1] We highlight that the (4) is exactly the minimum MSE (MMSE) estimator. An arbitrary non-optimal estimator except MMSE, which cannot apply the conditional expectation properties, may not satisfy the separable condition..
Theorem 1: The SDAS distortion $D=D_{s}+D_{c}$ is achievable within a separable distortion metric, if and only if
@@
where $C_{\text{IT}}(D_{s},B)$ and $R_{\text{IT}}(D_{c})$ are constrained channel capacity (7) and rate-distortion function (11), respectively.
@@ III-A Converse
We start with a converse to show that any achievable coding scheme must satisfy (14). Consider a $(2^{nR},n)$ coding scheme defined by the encoding and decoding functions $\phi$ and $\psi$. Let $\mathsf{\tilde{S}}^{n}=h^{\star}(\mathsf{X}^{n},\mathsf{Z}^{n})$ be the estimate sequence as given in (3) and $\mathsf{\hat{S}}^{n}=\psi(\phi(\mathsf{\tilde{S}}^{n}))$ be the reconstruction sequence.
@@ III-B Achievability
[…]
@@
$),(\mathsf{\tilde{S}}^{n},\mathsf{\hat{S}}^{n}(w))\in\mathcal{T}^{(n)}_{d,\epsilon_{s}}(P_{\mathsf{\tilde{S}\hat{S}}}(\tilde{s},\hat{s}$)), (17)
[…]
@@
[…]
@@
According to the achievability of lossy source coding [32, Theorem 10.2.1], we have $P_{e_{s}}$ tends to zero for sufficiently large $n$ whenever $R\geq R_{\text{IT}}(D_{c})$.
@@
[…]
according to the typical average lemma [20]. In summary, the SDAS distortion can be attained by
@@
where $(a)$ follows the condition of the separable distortion metric in MSST and we omit the terms containing the product of $P_{e_{s}}$ and $P_{e_{c}}$ in step $(b)$. Consequently, taking $n\to\infty$ and $P_{e_{s}},P_{e_{c}},\epsilon_{c},\epsilon_{e},\epsilon_{s}\to 0$, we can conclude that the expected SAS distortion (averaged over the random codebooks, state and channel noise) tends to be $D\to D_{c}+D_{s}$ whenever $R_{\text{IT}}(D_{c})\leq R\leq C_{\text{IT}}(D_{s},B)$.
@@
Remark: Intuitively, the sensing process determines the accuracy of the state information acquisition at the encoder, whereas the communication process governs the quantity of information transmitted to the decoder. Overemphasis on either process can result in substantial performance degradation of the SDAS system. Theorem 1 and its proof quantitatively show the condition (i.e., formula (14)) required to achieve the SDAS distortion $D$. Although the separable condition in (12) is to some extend strong, Theorem 1 is applicable in most parameter estimation scenarios which adopts MMSE estimator.
Additionally, Theorem 1 provides us a powerful tool to investigate the optimal ISAC channel input $\mathsf{X}$. Specifically, achieving optimal sensing performance alone may not ensure the accurate reconstruction at the decoder due to limitations in the communication channel capacity. Conversely, excessive communication capacity may be wasted if the sensory data lacks sufficient accuracy. This underscores the need to minimize the SDAS distortion by optimizing the distribution (Section IV) and the waveform (Section V) in next sections.
IV Channel Input Distribution Optimization
In this section, we explore the optimal input distribution for the SDAS system under scalar channels, aiming to minimize the SDAS distortion under the MMST constraint. Hereinafter, we adopt MSE as the distortion metric to guarantee the MSST.
@@ IV-A Problem Formulation
$\mathcal{P}_{0}\left\{\kern 2.0pt\begin{aligned} \mathop{\text{min}}\limits_{P_{\mathsf{X}}(x)}\kern 5.0pt&D=D_{c}+D_{s}\\ \text{subject to}\kern 5.0pt&R(D_{c})\leq I(D_{s},B),\end{aligned}\right$. (22)
where the definitions of $R(D_{c})$ and $I(D_{s},B)$ are given in (11) and (7), respectively. Here, we drop the subscript “IT” to emphasize the operational meanings is justified by the MSST from the perspective of coding scheme. Furthermore, we use the mutual information symbol $I(D_{s},B)$ instead of $C(D_{s},B)$ to emphasize that the optimal distribution of $\mathcal{P}_{0}$ does not necessarily achieve the channel capacity. This modification is just to keep the math rigorous, but it does not change the expression. By substituting the associated expressions into (22), problem $\mathcal{P}_{0}$ can be reformulated by
@@
[…]
@@
where the cost $e(x)$ defined in (6) can be expressed by
@@
[…]
@@
[…]
@@ IV-B Two-step BA-based Optimal Search Algorithm
2) Sub-problem 2: We calculate the minimum communication distortion $D_{c}$ for the given channel input distribution $P_{\mathsf{X}}(x)$ and determine the communication distortion $D_{c}$, thereby obtaining the SDAS distortion $D$.
By varying the preset sensing distortions, one may generate a set $\mathcal{D}$ collecting the SDAS distortions calculated through the above two steps. Thus, the minimum SDAS distortion and its corresponding distribution $P_{\mathsf{X}}(x)$ can be identified by the minimum value search. The detailed procedure is as follows.
$\bullet$ Constrained communication channel Capacity.
@@
$\mathop{\text{max}}\limits_{P_{\mathsf{X}}(x)}$ $I(\mathsf{X};\mathsf{Y})$ (30)
@@
$\mathop{\text{max}}\limits_{P_{\mathsf{X}}(x)}$ $I(P_{\mathsf{X}}(x),Q_{\mathsf{Y|X}}(y|x))-\mu\sum_{x}P_{\mathsf{X}}(x)e(x)$ (31)
@@
In formula (33), the Lagrangian multiplier $\lambda$ and the posterior distribution $Q_{\mathsf{X|Y}}(x|y)$ are unknown. The former may be determined by a bisection search method such that $\sum_{x}P_{\mathsf{X}}(x)b(x)\to B$. For the latter, we have the optimal posterior distribution which maximizes the MI with given $P_{\mathsf{X}}(x)$ and $Q_{\mathsf{Y|X}}(y|x)$, expressed by [33, Lemma 9.1]